# NASP: Network Slice as a Service Platform for 5G Networks

### v2

@@ Abstract
With the rapid global adoption of fifth-generation (5G) mobile telecommunications, the demand for highly flexible private networks has surged. A key beyond‑5G feature is network slicing, where the 3rd Generation Partnership Project (3GPP) defines three main use cases: massive Machine-Type Communications (mMTC), enhanced Mobile Broadband (eMBB), and Ultra-Reliable Low-Latency Communications (URLLC), along with their associated management functions. Similarly, the European Telecommunications Standards Institute (ETSI) provides the Zero-Touch Network and Service Management (ZSM) standard, enabling operation without human intervention. However, current technical documents lack definitions for end-to-end (E2E) management and integration across domains and subnet instances. We present a Network Slice as a Service Platform (NASP) that is agnostic to 3GPP and non-3GPP networks, addressing this gap. The NASP architecture comprises (i) onboarding requests for new slices at the business level, translating them into definitions of physical instances and interfaces among domains, (ii) a hierarchical orchestrator coordinating management functions, and (iii) communication interfaces with network controllers. Our NASP prototype is developed based on technical documents from 3GPP and ETSI, analyzing design overlaps and gaps across different perspectives. Results demonstrate the platform’s adaptability in handling diverse requests via the Communication Service Management Function. Evaluation indicates that Core configuration accounts for 68% of the time required to create a Network Slice Instance. Tests reveal a 93% reduction in data session establishment time when comparing URLLC and Shared scenarios. Finally, we present cost variations for operating the platform with the orchestration of five and ten slices, showing a 112% variation between Edge and Central deployments.
@@ 1 Introduction
The rapid, sustained increase in mobile data traffic, combined with the emerging diversity of service requirements from Massive Machine-Type Communication (mMTC) terminals and vertical enterprises, challenges the traditional mobile‐broadband paradigm (Navarro-Ortiz2020). Legacy networks, initially designed with only a limited set of parameters (e.g., priority class and Quality of Service (QoS) for broadband services), now face the task of accommodating heterogeneous and, in many cases, conflicting operational needs (Vila2020). For instance, a railway network demands seamless, high-speed mobility management along predefined routes, whereas an electricity metering system requires support for static, low-volume transmissions. Recent advances in Network Functions Virtualization (NFV) and Software-Defined Networking (SDN) have paved the way for innovations such as Network Slicing (NS), which enables the logical separation of network functions and resources tailored to specific technical and commercial requirements (Wijethilaka2021; Ksentini2020; Mahdi2023).
@@
NS has conceptual roots in earlier technologies, such as the 802.1Q virtual local‑area network standard (IEEE8021Q-2018) and Internet Engineering Task Force (IETF) RFC 4026 Virtual Private Networks (rfc4026), which provided isolated broadcast and session domains. However, deploying NS in cellular environments introduces additional challenges, including mobility management, control‑plane authentication, and session and charging management in the user plane (mahyoub2024security). Studies led by the Third Generation Partnership Project (3GPP) and initiatives by major industry players indicate an evolution toward End-to-End (E2E) network slices that integrate both 3GPP and non‑3GPP domains, enhancing connectivity for the Internet of Things (IoT) and enabling diverse service customization (Nguyen2016; bertenyi2018ng; Linnartz2022-lb). Despite these advancements, the complexity stemming from distributed interfaces across Radio Access Network (RAN), Transport Network (TN), and Core Network (CN) necessitates novel orchestration and management solutions (Wyszkowski2024; 7962822).
@@
The evolution toward heterogeneous mobile networks, driven by massive IoT deployments, emerging business-oriented services, and stringent performance requirements, motivates the need for a comprehensive, automated approach to NS. Existing studies have addressed resource allocation using machine‑learning techniques (Jiang; Li2018DeepRL) and have developed frameworks for NS control and assurance (Abbas2020IBNSlicingIN; Theodorou; Bega). However, fully integrated E2E orchestration that translates Service Level Agreement (SLA)/ Service Level Specification (SLS) requirements into practical NS designs remains an open challenge, as clear guidelines that bridge the gap between business-level intents and concrete slice configurations spanning the RAN, TN, and CN domains are still lacking in the literature.
@@
The main contributions of this work are threefold. First, we introduce NASP, an automated NSaaS platform that translates business‐rule templates into multi‐domain slice descriptors and orchestrates E2E deployments across both 3GPP and non‐3GPP networks. Second, we propose a hierarchical orchestration model, including Communication Service Management Function (CSMF), Network Slice Management Function (NSMF), and domain‐specific Network Slice Subnet Management Functions, which integrate closed‐loop quality assurance via CI/CD and IaC practices, enabling dynamic SLA compliance through AI‐driven telemetry analytics. Third, we develop and release a fully functional prototype built on Kubernetes, Open Network Operating System (ONOS), my5G-RANTester (Silveira2022), and Free 5G Core (Free5gc), demonstrating seamless integration, extensibility, and native support for physical and virtual resources in heterogeneous deployment environments.
Our comprehensive evaluation, covering mMTC, Ultra-Reliable Low Latency Communications (URLLC), shared, and non-3GPP use cases, yields several impactful findings. We observe that 66% of the E2E slice instantiation time is spent on Core domain configuration, pinpointing critical optimization targets. URLLC slices achieve up to a 93% reduction in data-session establishment time compared to shared slices. A prototype closed-loop quality assurance module demonstrates timely detection and mitigation of performance anomalies in our testbed. Finally, a cost-efficiency analysis reveals a 112% variation in monthly operational expenses between edge and centralized deployments, highlighting the economic differences between the two deployment options.
@@ 2 Background and Evolution of 5G Networks
The evolution of mobile telecommunications toward Fifth Generation (5G) introduced network programmability and NSaaS to better meet emerging service and market needs (3gpp.21.915; afolabi2018network). According to 3gpp.23.501, the 5G system and its Service-Based Architecture (SBA) were first specified in Releases 15–16, with continued refinements in later 3GPP releases (TS 23.501 v19.2.0, Rel-19). The architecture retains core elements, User Equipment (UE), Next Generation Radio Access Network (NG-RAN), and 5G Core (5GC), and separates control and user planes via Access and Mobility Management Function (AMF) and User Plane Function (UPF). Figure 1 highlights supporting functions such as Policy Control Function (PCF) (policy), Authentication Server Function (AUSF) (authentication), Network Slice Selection Function (NSSF) (slice selection), Network Data Analytics Function (NWDAF) (analytics), and Network Exposure Function (NEF) (exposure).
@@
NS is a central 5G innovation, allowing multiple logical networks to share a common infrastructure while meeting distinct SLAs. Each slice orchestrates resources across the RAN, TN, and CN and is characterized according to 3GPP and Global System for Mobile Communications Association (GSMA) guidance (3gpp.23.501; gsma-116) (GSMA NG.116 v10.0). The Network Slice Type (NEST) concept groups key attributes, throughput, latency, and reliability, into a service profile. For example, an Enhanced Mobile Broadband (eMBB) slice may target 99.999% availability, multimedia telephony, specific QoS, and strong session continuity. Moreover, isolation is essential since congestion or scaling in one slice must not degrade others, and separation must hold for performance, management, and security/privacy (3gpp.28.552). In this case, the system can be viewed in three domains, as can be seen in Figure 1: the RAN (gNB with Radio Unit (RU)/ Distributed Unit (DU)/ Centralized Unit (CU)), the TN (access–core transport), and the CN (interfacing with external Data Networks via the UPF).
Releases 17–18 (3gpp.21.917; 3gpp.21.918) extend slicing toward NSaaS, enabling operators to offer tenants customized, on-demand slices. NSaaS exposes management capabilities via standardized Application Programming Interfaces from 3GPP and TM Forum frameworks (gsma-116). Slice management relies on CSMF, NSMF, and NSSMF to translate tenant demands into an E2E deployment. Figure 2 outlines this orchestration, which includes an NS communication interface and a management domain comprising the NSMF, CSMF, and domain-specific entities (RAN-NSSMF, TN-NSSMF, 5GC-NSSMF). These coordinate instantiation, monitoring, and resource allocation across the NG-RAN, TN, and 5GC. Multiple slices (1–N) can coexist with dedicated resources and connect to external Data Networks through defined channels.
@@
NS lifecycle management, from preparation (blueprints, reservations) to operation (activation, monitoring, capacity planning) and decommissioning, benefits from modularity and isolation (5gcorebook-ns). The resulting operational complexity motivates Zero-touch Network and Service Management (ZSM), which applies automation, Machine Learning (ML), and Artificial Intelligence (AI) for self-configuration, self-monitoring, self-healing, and self-optimization (etsi-zsm). Therefore, beyond-5G systems build on these principles to deliver agile, customized services that meet diverse SLAs. Building on these advances, 3gpp.21.917; 3gpp.21.918 describe the latest 3GPP specifications (TS 28.541, TS 28.531, TS 28.622, TS 28.623; Releases 18–19), and together with etsi-zsm they outline a precise Network Resource Model (NRM) for standardized management of 5GC, NG-RAN, and NS. Function Information Object Classes, such as AMFFunction, SMFFunction, UPFFunction, GNBDUFunction, GNBCUCPFunction, encapsulates logical behaviors with well-defined attributes, states, and relationships. The subnetwork IOC provides a hierarchical container to organize ManagedElements into administrative domains. For slicing, NetworkSlice and NetworkSliceSubnet IOCs are augmented with ServiceProfile and SliceProfile data types that represent performance, coverage, capacity, and isolation requirements, aligning with GSMA Generic Network Slice Template (GST)/NG.116 service descriptors to maintain globally consistent SLA representations. These definitions are bound to solution sets in Extensible Markup Language (XML), JavaScript Object Notation (JSON), and YANG, supporting interoperability across multi-vendor environments and aligning with European Telecommunications Standards Institute (ETSI) ZSM principles of zero-touch automation.
Releases 18–19 (3gpp.21.918; 3gpp.21.919) also extend NS-management capabilities with admission-control policies, energy-efficiency indicators, and positioning/synchronization attributes, which support stronger SLA isolation and more sustainable operations. By codifying NRM objects, states, and notifications, these standards provide the foundation for E2E slice orchestration, dynamic adaptation, and exposure of NSaaS through defined open APIs. This framework remains consistent with the SBA vision of 5G and beyond while retaining flexibility for future enhancements. The instantiation of a Network Slice Instance (NSI) follows a standardized API sequence among Operations Support System (OSS)/ Business Support System (BSS), CSMF, NSMF, domain NSSMFs, NFVManagement and Orchestration (MANO), and the Inventory. Figure 3 illustrates the high-level flow and the interplay between GSMA, 3GPP/ETSI, and ETSI standards in this process. The process begins when the OSS/BSS captures a tenant request encoded with GSMA GST/NG.116, translating commercial requirements into a service order as specified by GSMA-NG116. The CSMF produces a ServiceProfile that maps high-level SLA attributes (throughput, latency, coverage, isolation) into a standards-compliant representation (3gpp.28.541).
@@
The NSMF interprets this profile and issues allocateNsi, which creates context for the new slice and initiates decomposition into domain-specific subnets (3gpp.28.531). Each NG-RAN, TN, and 5GC NSSMF receives allocateNssi with a SliceProfile that refines resource constraints, including spectrum allocations, transport bandwidth, and core-function placement rules (3gpp.28.531). The NSSMF performs design-time resolution to select building blocks, policies, and placement options, producing a resolved subnet design aligned with declarative operations in TS 28.531 (3gpp.28.531). It engages the NFVMANO stack to instantiate the required Virtual Network Functions, Cloud-native Network Functions, and network resources so that performance objectives are considered from the outset (ETSI-NFV-MANO). As resources become active, the NSSMF invokes createMOI to persist the corresponding Managed Object Instances in the Inventory, formalizing their existence and relationships (3gpp.28.622). The NSMF consolidates these MOIs into an E2E slice, linking subnets and updating state to reflect readiness.
Finally, it is essential to highlight that where adjustments are required, modifyMOIAttributes is used to fine-tune runtime parameters such as slice capacity, policy bindings, or service exposure (3gpp.28.623). In this way, the Inventory evolves into a live repository of the slice structure, supporting lifecycle tracking and auditing. Closed-loop assurance systems and exposure APIs allow incremental modifications (via modifyMOIAttributes) that help maintain SLA compliance across NG-RAN, TN, and 5GC domains (etsi-zsm). Binding GST $\rightarrow$ ServiceProfile $\rightarrow$ SliceProfile $\rightarrow$ runtime MOIs preserves business intent and provides traceability from intent to realization. In accordance with the ETSI ZSM framework described by etsi-zsm, these steps correspond to Automation Scopes AS1–AS4, encompassing (AS1) candidate evaluation, (AS2) candidate selection, (AS3) aggregation and configuration of slicing entities, and (AS4) generation of subnet templates and operational deployment. Coordinated use of allocateNsi, allocateNssi, createMOI, and modifyMOIAttributes across all actors supports repeatable, interoperable, and automatable slice creation in line with ETSI ZSM’s zero-touch management as detailed by etsi-zsm.
@@ 3 Related Work
ONAP Project L H H M M L
OSM Project L H H M M L
5GTours Project L H H M M L
5g-zorro Project M M M L L L
Baranda2020 Project L H H L L L
tranoris2021openslice Project M H H M M L
camara2025 Project M H H M
opencapif2025 Project M H M M L L
nephio2025 Project H H H H M M
sylva2024 Project M H H M L L
hexaX2 Project H M H H M M
monb5g Project H M M H M L
daemon Project H M H H M L
Li2018DeepRL Article L H L L L L
Jiang Article M H L H M L
Bega Article M H L L H L
Abbas2020IBNSlicingIN Article M H M H M L
Fernandez Article M H M M L L
Theodorou Article H H L H H L
Larrea2023 Article M H L L M L
Scotece2023-5gkube Article M M M H L L
mlopez2023mapek Article H L M H M L
Wyszkowski2024 Article M L H L L L
Dalgitsis2024-kp Article H H M M H L
Esmat2024 Article H H H M H L
Chowdhury2024 Article H H H H M L
mlops2024 Article H L H H H M
zhao2025 Article H H H H M L
NASP (This Work) Article H H H H H H
@@
Several open-source initiatives automate 5G slice orchestration but vary in scope. ONAP and OSM offer comprehensive MANO frameworks but require significant integration effort for full E2E automation. Building on these, subsequent projects such as 5GTours, 5g-zorro, and Baranda2020 demonstrated vertical use cases but reported limited architectural evolution. Since 2023, the landscape has shifted toward B5G and AI-native solutions. Large-scale research programmes such as hexaX2 investigate B5G architectures that emphasize AI-driven automation and sustainability. Projects such as monb5g and daemon develop AI/ML-powered monitoring and control loops, aligning with the ETSI ZSM vision. The MAPE-K framework introduced by mlopez2023mapek is now a widely cited pattern for self-adaptive systems. Moreover, the MLOps practices cataloged by mlops2024 are being integrated into network automation to streamline the lifecycle of AI models in production networks. Furthermore, community-driven initiatives such as camara2025, opencapif2025, nephio2025, and sylva2024 develop open, cloud-native API and automation frameworks backed by major operators and vendors.
@@
Pushing lifecycle automation further, OpenSlice, as described by tranoris2021openslice, adopts TM Forum’s Open Digital Architecture to drive catalog-based order decomposition that supports zero-touch onboarding, scaling, and retirement of NFV artifacts. However, it delegates granular 5GC tasks, such as UPF placement, PCF/ Network Repository Function (NRF) affinity, and network-exposure APIs, to external MANO stacks, excluding the SBA from the reconfiguration loop. In response, the NASP framework presented in this article unifies hierarchical slice composition, AI-powered assurance, and tenant-facing NSaaS exposure across RAN, TN, and CN domains.
@@
AI-driven research is steadily closing specific gaps in the NS lifecycle. Li2018DeepRL were the first to apply Deep Reinforcement Learning (DRL) for joint radio- and core-resource balancing under fluctuating loads. Jiang introduced intelligence slicing, a unified framework that instantiates AI modules, such as Recurrent Neural Network (RNN)-based channel prediction and security-anomaly detection, on demand. Bega integrated a 3-D Convolution Neural Network (CNN) capacity-forecasting model (DeepCog) with Reinforcement Learning (RL) control loops, enabling proactive VNF scaling throughout the slice lifecycle. Abbas2020IBNSlicingIN proposed IBNSlicing, an intent-based, E2E framework that automatically translates high-level slice “contracts” into Open Source MANO (OSM) and FlexRAN templates; a hybrid Generative Adversarial Network (GAN), i.e., Long Short-term Memory (LSTM) + CNN, forecasts slice-level CPU and RAM requirements to trigger proactive scaling and admission decisions. Moreover, Fernandez designed a Distributed Ledger Technologies (DLT)-enabled 5G marketplace in which multiple providers securely advertise, negotiate, and trade heterogeneous resources via smart-contract automation, validating the concept with a Corda-based VNF-trading proof of concept. Complementing these point solutions, Wyszkowski2024 proposed a taxonomy of network-slice-instance patterns that now guides subsequent work.
@@
Extending the state of the art, Theodorou outlined a blockchain-based, zero-touch service-assurance loop for cross-domain slicing, in which smart-contract-bound SLAs feed AI-driven monitoring, prediction, and automated reconfiguration across the lifecycle. Dalgitsis2024-kp proposed a cloud-native Slice Federation-as-a-Service (SFaaS) framework that augments the GSMA Operator Platform with a new east–west APIs, enabling seamless slice federation across operators and demonstrating the approach on a live 5G/B5G testbed. Esmat2024 decomposed carrier-grade SLAs into federated slice objectives, whereas Chowdhury2024 introduced Accelerator, an intent-driven, zero-touch resource-slicing algorithm that places and schedules Service Function Chain (SFC) workloads for B5G and legacy services across SDN/NFV infrastructures.

### v3

@@ Abstract
With the rapid global adoption of fifth-generation (5G) mobile telecommunications, the demand for highly flexible private networks has surged. A key beyond‑5G feature is network slicing, where the 3rd Generation Partnership Project (3GPP) defines three main use cases: massive Machine-Type Communications (mMTC), enhanced Mobile Broadband (eMBB), and Ultra-Reliable Low-Latency Communications (URLLC), along with their associated management functions. Similarly, the European Telecommunications Standards Institute (ETSI) provides the Zero-Touch Network and Service Management (ZSM) standard, enabling operation without human intervention. However, current technical documents lack definitions for end-to-end (E2E) management and integration across domains and subnet instances. We present a Network Slice as a Service Platform (NASP) that is agnostic to 3GPP and non-3GPP networks, addressing this gap. The NASP architecture comprises (i) onboarding requests for new slices at the business level, translating them into definitions of physical instances and interfaces among domains, (ii) a hierarchical orchestrator coordinating management functions, and (iii) communication interfaces with network controllers. Our NASP prototype is developed based on technical documents from 3GPP and ETSI, analyzing design overlaps and gaps across different perspectives. Results demonstrate the platform’s adaptability in handling diverse requests via the Communication Service Management Function. Evaluation indicates that the Core configuration accounts for 68% of the time required to create a Network Slice Instance. Tests reveal a 93% reduction in data session establishment time when comparing URLLC and Shared scenarios, i.e., an eMBB-type slice that reuses existing control-plane Network Functions to represent a best-effort provisioning baseline. Finally, we present cost variations for operating the platform with the orchestration of five and ten slices, showing a 112% variation between Edge and Central deployments.
@@ 1 Introduction
The rapid, sustained increase in mobile data traffic, combined with the emerging diversity of service requirements from Massive Machine-Type Communication (mMTC) terminals and vertical enterprises, challenges the traditional mobile‐broadband paradigm (Navarro-Ortiz and others, 2020). Legacy networks, initially designed with only a limited set of parameters (e.g., priority class and Quality of Service (QoS) for broadband services), now face the task of accommodating heterogeneous and, in many cases, conflicting operational needs (Vilà and others, 2020). For instance, a railway network demands seamless, high-speed mobility management along predefined routes, whereas an electricity metering system requires support for static, low-volume transmissions. Recent advances in Network Functions Virtualization (NFV) and Software-Defined Networking (SDN) have paved the way for innovations such as Network Slicing (NS), which enables the logical separation of network functions and resources tailored to specific technical and commercial requirements (Wijethilaka and Liyanage, 2021; Ksentini and Frangoudis, 2020; Mahdi and Abdullah, 2023).
@@
NS has conceptual roots in earlier technologies, such as the 802.1Q virtual local‑area network standard (IEEE, 2018) and Internet Engineering Task Force (IETF) RFC 4026 Virtual Private Networks (Andersson and Madsen, 2005), which provided isolated broadcast and session domains. However, deploying NS in cellular environments introduces additional challenges, including mobility management, control‑plane authentication, and session and charging management in the user plane (Mahyoub and others, 2024). Studies led by the Third Generation Partnership Project (3GPP) and initiatives by major industry players indicate an evolution toward End-to-End (E2E) network slices that integrate both 3GPP and non‑3GPP domains, enhancing connectivity for the Internet of Things (IoT) and enabling diverse service customization (Nguyen and others, 2016; Bertenyi and others, 2018; Linnartz and others, 2022). Despite these advancements, the complexity stemming from distributed interfaces across Radio Access Network (RAN), Transport Network (TN), and Core Network (CN) necessitates novel orchestration and management solutions (Wyszkowski and others, 2024; Devlic and others, 2017).
@@
The evolution toward heterogeneous mobile networks, driven by massive IoT deployments, emerging business-oriented services, and stringent performance requirements, motivates the need for a comprehensive, automated approach to NS. Existing studies have addressed resource allocation using machine‑learning techniques (Jiang et al., 2019; Li et al., 2018) and have developed frameworks for NS control and assurance (Abbas and others, 2020; Theodorou and others, 2021; Bega and others, 2020). However, fully integrated E2E orchestration that translates Service Level Agreement (SLA)/ Service Level Specification (SLS) requirements into practical NS designs remains an open challenge, as clear guidelines that bridge the gap between business-level intents and concrete slice configurations spanning the RAN, TN, and CN domains are still lacking in the literature.
@@
The main contributions of this work are threefold. First, we introduce NASP, an automated NSaaS platform that translates business‐rule templates into multi‐domain slice descriptors and orchestrates E2E deployments across both 3GPP and non‐3GPP networks. Second, we propose a hierarchical orchestration model, including Communication Service Management Function (CSMF), Network Slice Management Function (NSMF), and domain‐specific Network Slice Subnet Management Functions, which integrate closed‐loop quality assurance via CI/CD and IaC practices, enabling dynamic SLA compliance through AI‐driven telemetry analytics. Third, we develop and release a fully functional prototype built on Kubernetes, Open Network Operating System (ONOS), my5G-RANTester (Silveira and others, 2022), and Free 5G Core (Free5gc), demonstrating seamless integration, extensibility, and native support for physical and virtual resources in heterogeneous deployment environments.
Our comprehensive evaluation, covering mMTC, Ultra-Reliable Low Latency Communications (URLLC), shared, and non-3GPP use cases, yields several impactful findings. In this context, the shared use case corresponds to an Enhanced Mobile Broadband (eMBB)-type slice that reuses existing control-plane Network Functions, serving as a best-effort baseline for comparison against fully isolated deployments. We observe that 66% of the E2E slice instantiation time is spent on Core domain configuration, pinpointing critical optimization targets. URLLC slices achieve up to a 93% reduction in data-session establishment time compared to shared slices. A prototype closed-loop quality assurance module demonstrates timely detection and mitigation of performance anomalies in our testbed. Finally, a cost-efficiency analysis reveals a 112% variation in monthly operational expenses between edge and centralized deployments, highlighting the economic differences between the two deployment options.
@@ 2 Background and Evolution of 5G Networks
The evolution of mobile telecommunications toward Fifth Generation (5G) introduced network programmability and NSaaS to better meet emerging service and market needs (3GPP, 2019a; Afolabi and others, 2018). According to 3GPP (2025e), the 5G system and its Service-Based Architecture (SBA) were first specified in Releases 15–16, with continued refinements in later 3GPP releases (TS 23.501 v19.2.0, Rel-19). The architecture retains core elements, User Equipment (UE), Next Generation Radio Access Network (NG-RAN), and 5G Core (5GC), and separates control and user planes via Access and Mobility Management Function (AMF) and User Plane Function (UPF). Figure 1 highlights supporting functions such as Policy Control Function (PCF) (policy), Authentication Server Function (AUSF) (authentication), Network Slice Selection Function (NSSF) (slice selection), Network Data Analytics Function (NWDAF) (analytics), and Network Exposure Function (NEF) (exposure).
@@
NS is a central 5G innovation, allowing multiple logical networks to share a common infrastructure while meeting distinct SLAs. Each slice orchestrates resources across the RAN, TN, and CN and is characterized according to 3GPP and Global System for Mobile Communications Association (GSMA) guidance (3GPP, 2025e; GSMA, 2025) (GSMA NG.116 v10.0). The Network Slice Type (NEST) concept groups key attributes, throughput, latency, and reliability, into a service profile. For example, an eMBB slice may target 99.999% availability, multimedia telephony, specific QoS, and strong session continuity. Moreover, isolation is essential since congestion or scaling in one slice must not degrade others, and separation must hold for performance, management, and security/privacy (3GPP, 2025d). In this case, the system can be viewed in three domains, as can be seen in Figure 1: the RAN (gNB with Radio Unit (RU)/ Distributed Unit (DU)/ Centralized Unit (CU)), the TN (access–core transport), and the CN (interfacing with external Data Networks via the UPF).
Releases 17–18 (3GPP, 2022b, 2024c) extend slicing toward NSaaS, enabling operators to offer tenants customized, on-demand slices. NSaaS exposes management capabilities via standardized Application Programming Interfaces from 3GPP and TM Forum frameworks (GSMA, 2025). Slice management relies on CSMF, NSMF, and NSSMF to translate tenant demands into an E2E deployment. Figure 2 outlines this orchestration, which includes an NS communication interface and a management domain comprising the NSMF, CSMF, and domain-specific entities (RAN-NSSMF, TN-NSSMF, 5GC-NSSMF). These coordinate instantiation, monitoring, and resource allocation across the NG-RAN, TN, and 5GC. Multiple slices (1–N) can coexist with dedicated resources and connect to external Data Networks through defined channels.
@@
NS lifecycle management, from preparation (blueprints, reservations) to operation (activation, monitoring, capacity planning) and decommissioning, benefits from modularity and isolation (Rommer and others, 2020). The resulting operational complexity motivates Zero-touch Network and Service Management (ZSM), which applies automation, Machine Learning (ML), and Artificial Intelligence (AI) for self-configuration, self-monitoring, self-healing, and self-optimization (ETSI, 2024). Therefore, beyond-5G systems build on these principles to deliver agile, customized services that meet diverse SLAs. Building on these advances, the latest 3GPP specifications (Rel-18/19) outline a precise Network Resource Model (NRM) utilizing Information Object Classes (e.g., NetworkSlice, NetworkSliceSubnet) to standardize management (3GPP, 2024c). These are augmented with ServiceProfile and SliceProfile descriptors that align with GSMA Generic Network Slice Template (GST) to maintain globally consistent SLA representations. By codifying NRM objects, states, and operations across multi-vendor definitions ( Extensible Markup Language (XML), JavaScript Object Notation (JSON), YANG), these standards provide the necessary foundation for E2E slice orchestration and exposure of NSaaS through open APIs. The instantiation of a Network Slice Instance (NSI) follows a standardized API sequence among Operations Support System (OSS)/ Business Support System (BSS), CSMF, NSMF, domain NSSMFs, NFVManagement and Orchestration (MANO), and the Inventory. Figure 3 illustrates the high-level flow and the interplay between GSMA, 3GPP/ European Telecommunications Standards Institute (ETSI), and ETSI standards in this process. The process begins when the OSS/BSS captures a tenant request encoded with GSMA GST/NG.116, translating commercial requirements into a service order as specified by GSMA (2023). The CSMF produces a ServiceProfile that maps high-level SLA attributes (throughput, latency, coverage, isolation) into a standards-compliant representation (3GPP, 2025a).
@@
Following GST modeling, the NSMF translates the ServiceProfile into an allocateNsi request to decompose the slice into domain-specific subnets. Each domain NSSMF resolves the SliceProfile and engages the NFVMANO stack to instantiate resources, which are persisted as Managed Object Instances via createMOI 3GPP (2025b); ETSI (2021). When dynamic adjustments are necessary, runtime parameter tuning is achieved through modifyMOIAttributes (3GPP, 2025c). This structured progression from business intent (GST $\rightarrow$ ServiceProfile $\rightarrow$ SliceProfile) to runtime MOIs corresponds directly to the ETSI ZSM Automation Scopes AS1–AS4 (ETSI, 2024), enabling a repeatable, zero-touch orchestration process.
@@ 3 Related Work
ONAP (2022) Project L H H M M L
ETSI MANO (2022) Project L H H M M L
5G Tours (2024) Project L H H M M L
Breitgand and others (2021) Project M M M L L L
Baranda and others (2020) Project L H H L L L
Tranoris (2021) Project M H H M M L
CAMARA Project Consortium (2025) Project M H H M
ETSI TeraFlowSDN and OpenCAPIF Team (2025) Project M H M M L L
Nephio Project Community (2025) Project H H H H M M
Project Sylva Consortium (2024) Project M H H M L L
Hexa-X-II Consortium / SNS JU (2025) Project H M H H M M
MonB5G Project Consortium (2023) Project H M M H M L
DAEMON Project Consortium (2023) Project H M H H M L
Li et al. (2018) Article L H L L L L
Jiang et al. (2019) Article M H L H M L
Bega and others (2020) Article M H L L H L
Abbas and others (2020) Article M H M H M L
Fernández-Fernández and others (2021) Article M H M M L L
Theodorou and others (2021) Article H H L H H L
Larrea et al. (2023) Article M H L L M L
Scotece and others (2023) Article M M M H L L
López and others (2023) Article H L M H M L
Wyszkowski and others (2024) Article M L H L L L
Dalgitsis and others (2024) Article H H M M H L
Esmat and Lorenzo (2024) Article H H H M H L
Chowdhury (2024) Article H H H H M L
Smith and Doe (2024) Article H L H H H M
Zhao et al. (2025) Article H H H H M L
NASP (This Work) Article M H H H H H
@@
Several open-source initiatives automate 5G slice orchestration but vary in scope. ONAP (2022) and ETSI MANO (2022) offer comprehensive MANO frameworks but require significant integration effort for full E2E automation. Building on these, subsequent projects such as 5G Tours (2024), Breitgand and others (2021), and Baranda and others (2020) demonstrated vertical use cases but reported limited architectural evolution. Since 2023, the landscape has shifted toward B5G and AI-native solutions. Large-scale research programmes such as Hexa-X-II Consortium / SNS JU (2025) investigate B5G architectures that emphasize AI-driven automation and sustainability. Projects such as MonB5G Project Consortium (2023) and DAEMON Project Consortium (2023) develop AI/ML-powered monitoring and control loops, aligning with the ETSI ZSM vision. The MAPE-K framework introduced by López and others (2023) is now a widely cited pattern for self-adaptive systems. Moreover, the MLOps practices cataloged by Smith and Doe (2024) are being integrated into network automation to streamline the lifecycle of AI models in production networks. Furthermore, community-driven initiatives such as CAMARA Project Consortium (2025), ETSI TeraFlowSDN and OpenCAPIF Team (2025), Nephio Project Community (2025), and Project Sylva Consortium (2024) develop open, cloud-native API and automation frameworks backed by major operators and vendors.
@@
Pushing lifecycle automation further, OpenSlice, as described by Tranoris (2021), adopts TM Forum’s Open Digital Architecture to drive catalog-based order decomposition that supports zero-touch onboarding, scaling, and retirement of NFV artifacts. However, it delegates granular 5GC tasks, such as UPF placement, PCF/ Network Repository Function (NRF) affinity, and network-exposure APIs, to external MANO stacks, excluding the SBA from the reconfiguration loop. In response, the NASP framework presented in this article unifies hierarchical slice composition, AI-powered assurance, and tenant-facing NSaaS exposure across RAN, TN, and CN domains.
@@
AI-driven research is steadily closing specific gaps in the NS lifecycle. Li et al. (2018) were the first to apply Deep Reinforcement Learning (DRL) for joint radio- and core-resource balancing under fluctuating loads. Jiang et al. (2019) introduced intelligence slicing, a unified framework that instantiates AI modules, such as Recurrent Neural Network (RNN)-based channel prediction and security-anomaly detection, on demand. Bega and others (2020) integrated a 3-D Convolution Neural Network (CNN) capacity-forecasting model (DeepCog) with Reinforcement Learning (RL) control loops, enabling proactive Virtual Network Function (VNF) scaling throughout the slice lifecycle. Abbas and others (2020) proposed IBNSlicing, an intent-based, E2E framework that automatically translates high-level slice “contracts” into Open Source MANO (OSM) and FlexRAN templates; a hybrid Generative Adversarial Network (GAN), i.e., Long Short-term Memory (LSTM) + CNN, forecasts slice-level CPU and RAM requirements to trigger proactive scaling and admission decisions. Moreover, Fernández-Fernández and others (2021) designed a Distributed Ledger Technologies (DLT)-enabled 5G marketplace in which multiple providers securely advertise, negotiate, and trade heterogeneous resources via smart-contract automation, validating the concept with a Corda-based VNF-trading proof of concept. Complementing these point solutions, Wyszkowski and others (2024) proposed a taxonomy of network-slice-instance patterns that now guides subsequent work.
@@
Extending the state of the art, Theodorou and others (2021) outlined a blockchain-based, zero-touch service-assurance loop for cross-domain slicing, in which smart-contract-bound SLAs feed AI-driven monitoring, prediction, and automated reconfiguration across the lifecycle. Dalgitsis and others (2024) proposed a cloud-native Slice Federation-as-a-Service (SFaaS) framework that augments the GSMA Operator Platform with a new east–west APIs, enabling seamless slice federation across operators and demonstrating the approach on a live 5G/B5G testbed. Esmat and Lorenzo (2024) decomposed carrier-grade SLAs into federated slice objectives, whereas Chowdhury (2024) introduced Accelerator, an intent-driven, zero-touch resource-slicing algorithm that places and schedules Service Function Chain (SFC) workloads for B5G and legacy services across SDN/NFV infrastructures.